OpenStudy (anonymous):

write the slope-intercept form of the equation of the line described: through: (2,-1) parallel to y=-2x+2

OpenStudy (anonymous):

i know that the slope will be the same, since it's parallel.. i just am not sure as to how to find the y-intercept

OpenStudy (anonymous):

\[y-(-1)=-2(x-2)\]\[y=-2x+4-1\]\[y=-2x+3\]

OpenStudy (anonymous):

is the formula: y-y1=M(x-x1)?

OpenStudy (anonymous):

???

OpenStudy (noseboy908):

for point-intercept, yes. it is.

OpenStudy (anonymous):

okay, thanks. you mind helping me with perpindicular real fast?

OpenStudy (noseboy908):

sure.

OpenStudy (anonymous):

okay thanks. it says: through: (-5,2) perp. to y=-5/3x+1

OpenStudy (anonymous):

anyone??

OpenStudy (noseboy908):

perp means that it's slope is the negative rec. of your other given slope, so plug in the new point and slope, and solve.

OpenStudy (anonymous):

so my slope is now 3/5?

OpenStudy (noseboy908):

yep.

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

is 3/5x+1 correct?

OpenStudy (noseboy908):

no, the slope is correct, but the y-int isn't.

OpenStudy (anonymous):

okay i seee my mistake

OpenStudy (anonymous):

y=3/5x+4 ?

OpenStudy (noseboy908):

still nope

OpenStudy (anonymous):

what's the answer?